# Spike-frequency and h-current based adaptation are dynamically equivalent in a Wilson-Cowan field model

### v2

@@ IV Characterization of regular spatio-temporal patterns in Hopf- and Turing-unstable states
Figure 5: Properties of the spatio-temporal patterns emerging in Hopf- and Turing-unstable states for increasing h-current strength. A $k_{max}$ values (top row), dominant spatial frequency $f_{x}$ (middle row), and dominant temporal frequency $f_{t}$ (bottom row) for each location in state space. The figure shows slices of state space spanned by the external input $(I_{e},I_{e})\in[-2.5,2]\times[-3.5,1]$ for $b\in\{0,-0.5,-1.0\}$. Purple colors denote the corresponding feature values. Darker colors indicate higher values, see color bar. Yellow denotes states of no spatio-temporal activity, light green denotes states with irregular patterns with a standard deviation $r$ of the Kuramoto-order parameter of $r>10^{-2}$. Thin lines denote the boundaries of the dynamical regimes shown in Figure 2A. Green boxes correspond to the panels shown in D. B Mean (dots) and variance (bars) of $k_{max},\ k_{0},\ f_{t},$ and $f_{x}$ for $b\in\{-0.25,-0.5,-0.75,-1\}$ for the states located in the Hopf-unstable, Turing-unstable up-, and Turing-unstable down-state regimes, for regular spatio-temporal activity for which the standard deviation of the Kuramoto-order parameter over time and space was $r\leq 10^{-2}$ (see Appendix B.3). C Histograms and kernel density estimates (KDE; thin lines) of the numerically calculated spatial frequency ($f_{x}$; top row), the spatial frequency corresponding to $k_{0}$ ($\frac{L\cdot k_{0}}{2\pi}$; middle row), and the spatial frequency corresponding to $k_{max}$ ($\frac{L\cdot k_{max}}{2\pi}$; bottom row) in the regimes of Hopf instability (left column), Turing instability in down-states (middle column), and Turing instability in up-states (right column) for b=0.5b=-0.5. Dashed vertical lines with numbers denote the location of the peak of the KDE. D Close-ups of state space (see green boxes in A) with values for $k_{max}$ (lower right), $f_{x}$ (upper right), and $f_{t}$ (upper left) in the Turing-unstable up- (left panels) and down-states (right panels) in bi. Yellow denotes states of no spatio-temporal activity or with irregular patterns with a standard deviation $r$ of the Kuramoto-order parameter of $r>10^{-2}$. The other colors denote feature values (see color bars). Wavenumbers were acquired as described in Section II.2; spatial and temporal frequencies as described in the Appendices B.1 and B.2. Numerical simulations were conducted with initialization close to the corresponding unstable fixed points.
@@
Figure 5 shows the results of the numerical investigation for the system with h-currents with increasing adaptation strength for the regimes of Hopf or Turing instability, where regular patterns occur. We did not consider the regimes of three fixed points (corresponding to dark and light beige in Figure 2A), because numerical simulations close to the fixed points undergoing a Turing bifurcation lead to irregular propagation of activity fronts similar to the pattern exemplified in Figure 2B, panel 7, which shows irregular activity emerging in a Turing-unstable up-state in bi.
@@
Figure 5A shows the wavenumber $k_{max}$ (top row), and the spatial, $f_{x}$ (center row) and temporal frequencies, $f_{t}$ (bottom row) of the emerging patterns for different locations in state space and for increasing adaptation strength. $k_{max}$ values vary continuously across state space. Values are larger than $0.44$ with highest values occurring in Turing-unstable states (see color bar). The spatial frequency, $f_{x}$, of the wave patterns change smoothly across state space for $b=0$ with the highest values occurring toward the center of the Hopf regime. Values decrease toward the borders. For b=0.5b=-0.5 and b=1b=-1, the distribution of $f_{x}$ becomes less smooth with non-equidistant, irregular patterns ($r>10^{-2}$) emerging toward the borders of the Hopf regime, where the regimes of Turing instability, one stable fixed point, Hopf instability, and three fixed points with only one stable state meet (green areas inside instability regimes, Figure 5A). The numerical simulations also confirm that the activity traces for $b=0$ are stationary spatial waves, as $f_{t}$ is zero. For adaptation strong enough, waves start moving resulting in a finite value of $f_{t}$ which increases with the value of $|b|$ (see also Figure 5B) confirming what is expected from the mathematical analysis. For increasing $|b|$, the averages of the wavenumbers $k_{0}$, and $k_{max}$ increase in value, albeit the variance over state space is high. $f_{x}$ increases with $|b|$ consistent with the increase of the wavenumbers $k_{0}$ and $k_{max}$. Additionally, the average temporal frequency increases steeply, which is expected from the increasing imaginary parts at the zero crossings of the real parts of the eigenvalues as discussed in Section III.
In Figure 5C we compare the spatial frequencies of the simulated activity patterns (top row) with the expected spatial frequencies corresponding to the two wavenumbers $k_{0}$, i.e., $\frac{L\cdot k_{0}}{2\pi}$ (middle row), and $k_{max}$, i.e., $\frac{L\cdot k_{max}}{2\pi}$ (bottom row). Since $k_{max}$ identifies the most rapidly growing unstable mode, expectations are to have highest similarity of spatial frequencies and frequencies corresponding to $k_{max}$. The distribution of spatial frequencies across Hopf states roughly agrees with this hypothesis (The Wasserstein distance between the distributions of $f_{x}$ and $\frac{L\cdot k_{max}}{2\pi}$ is $3.2$ and $12.9$ between the distributions of $f_{x}$ and $\frac{L\cdot k_{0}}{2\pi}$.), the spatial frequencies observed in Turing-unstable down- and up-states, however, agree better with $k_{0}$ (The Wasserstein distance between the distributions of $f_{x}$ and $\frac{L\cdot k_{max}}{2\pi}$ are larger than $14$ and smaller than $10$ between the distributions of $f_{x}$ and $\frac{L\cdot k_{0}}{2\pi}$.).
Moreover, spatial frequencies in Hopf-unstable states often lie in between $k_{0}$ and $k_{max}$, whereas for Turing-unstable states values are slightly below the spatial frequencies corresponding to $k_{0}$ which likewise held for simulated example activity traces in Hopf- and Turing-unstable states. These findings confirm that there is a range of spatial modes whose properties the activity patterns can inherit (called stability balloon in [63] [^5] In [63], a stability balloon describes a region in an $(R,\vec{q})$ plane, where $R$ is a control parameter to take the system further away from the equilibrium and $\vec{q}$ the wave vector effectively representing the wavenumber in a spatially one-dimensional ring.). Since we summarize fixed points stable at $k=0$ undergoing a Turing bifurcation at $k_{0}>0$ under Turing-unstable states, we know that $k_{0}$ is the smallest wavenumber at which the homogeneous equilibrium becomes unstable, i.e., determining the broadest wave. Nonetheless, we see that waves tend to converge to spatio-temporal patterns with a spatial frequency slightly below the expected frequency corresponding to the smallest wavenumber $k_{0}$, all the while the exact frequency varies depending on the initialization (see Appendix B.7, Figure 11).
Since we observe bistability in Hopf-unstable states between spatially homogeneous oscillations and temporally stationary spatial oscillations (see Section III, we performed the same analysis as in Figures 5A and D in Appendix B.5 but initialized close to zero. Results are similar and shown in Appendix B.5 Figure 10. This figure also shows the $k_{0}$ values over slices of state space corresponding to Figure 5A.
Figure 5D shows close-ups of Turing unstable up- and down-states in bi for b=0.5b=-0.5. $k_{max}$ values do not deviate strongly from the $k_{max}$ values of the rest of the considered state spaces but only few numerical simulations converged to regular spatio-temporal patterns (non-yellow locations in the close-ups of state space). The resulting patterns strongly depend on the initialization, and even initializations close to the Turing-unstable fixed point do not necessarily converge to a spatio-temporal pattern but to the coexisting stable fixed point, as exemplified in Figure 2B, panel 5(*), or irregular patterns [^6] States marked in yellow in the regions of Turing-unstable up- and down-states in bi shown in Figure 5D and Figure 6E may also support regular spatio-temporal patterns if another initialization would have been chosen., as exemplified in Figure 2B, panel 7. If regular spatio-temporal patterns emerge, traveling waves show low values of spatial frequency. Temporal frequencies are also lower compared to the rest of state space.
@@
Figure 6: Propagation speed of spatio-temporal patterns emerging in Hopf- and Turing-unstable states for increasing adaptation strength. A Slices of state space spanned by the external input currents showing speed values for the corresponding location in state space. Adaptation strength increases from left to right. Speed values are indicated by purple (see color bar), yellow denotes states of no spatio-temporal activity, green denotes states of irregular patterns with $r>10^{-2}$. Thin lines denote the borders of the dynamical regimes shown in Figure 2A. B Histograms of the distribution of speed values across the states shown in A. Dashed vertical lines and numbers denote the location of the peaks of the histograms. C Average (dots) and variance (bars) of speed values across state space for $b\in\{-0.25,-0.5,-0.75,-1.0\}$. D spatio-temporal patterns of $u_{e}(x,t)$ for different values of $b$ emerging in the locations in state space which are indicated by hexagons in A. The corresponding value of the external input currents are given in Table 4 in Appendix B.6. The corresponding speed values are for 1: <<0.5; 2 and 3: 11.6; 4: 17.7; 5: 14.6; 6: 18.7. All other parameters are given in Table 1. E Close-ups of state spaces (see green boxes in Figure 5A) with values for speed. Numerical simulations were initialized around the fixed points of Hopf or Turing instability. Speed values were computed as described in Appendix B.4.
@@
Figure 6B shows the distribution of speed values across Hopf-unstable and Turing-unstable states. The stronger the adaptation, the higher the speed values are. For $b=0$, as expected, activity patterns do not travel across space (recall Figure 2, panel 0) or their speed is <0.5<0.5 (see Figure 6D, panel 1). For b=0.5b=-0.5, there are two different speed values appearing the most often, with the dominant speed being $11.2$. For b=1b=-1, the variance of the distribution is larger and the most common speed value is once again higher at $17.6$. Figure 6C confirms that the average speed across state space increases with adaptation strength. Figure 6D shows traveling waves for the same relative locations in state space (i.e., considering the shift to the left discussed in Section III), increasing in adaptation strength from left to right column. “Broad" patterns become “narrower", increasing both temporal and spatial frequencies and their slope in the space-time diagrams, henceforth, the speed.
Figure 6E shows the same close-ups as Figure 5D of Turing-unstable up- and down-states in bi showing the speed values for the numerical simulations converging to regular traveling waves. Speed values tend to be higher closer to the bifurcation line to the regime of single Turing-unstable up- and down-states.
@@ V Discussion
A semi-analytical stability analysis across parameter space was performed for increasing adaptation strength resulting in the identification of over 10 dynamical regimes including two types of bistability. Spatio-temporal patterns, most often periodic adaptation-driven traveling waves, emerge in states exhibiting Hopf and Turing instabilities if adaptation is sufficiently strong. Activity patterns remain stationary in the absence of adaptation which is explained by the zero imaginary part of all eigenvalues of the linearization matrix when a fixed point undergoes a Turing bifurcation. This aligns with the results presented in [43], who prove the existence of stationary patterns for the reduced system if adaptation is weak enough while traveling waves occur for stronger adaptation.
Changing adaptation strength does not only alter the model’s dynamical regimes but also the features of the emerging temporal and spatial patterns including the speed of traveling wave activity. The numerical simulations aligned with the mathematical analysis performed before. With stronger adaptation, wavenumbers increase on average, causing more waves on the ring, i.e., higher dominant spatial frequencies. Analogously, with growing imaginary parts of the eigenvalues at the zero crossing of their respective real parts for stronger adaptation, temporal frequencies also increase. These changes correlate with increasing traveling speeds of activity patterns when the adaptation becomes stronger.
@@
In our population-based model, both adaptation mechanisms were implemented in a simplified way. Hyperpolarising feedback current of somatic SFA can be mediated by potassium channels, which activate when the excitatory neurons are in a depolarized high-activity state, effectively causing K+ outward flow [11, 65]. Different potassium channels have different properties leading to differential effects in the actual adaptation mechanism: Voltage- ($I_{m}$, see [66]) and sodium-activated K+-channels ($I_{KNa}$, see [67]) contribute mostly to subtreshold, while Ca2+-activated K+-channels ($I_{KCa}$, [68]) contribute to spike-based adaptation. Different channel types also have different effects on the response properties of a neuron, for example, on a neurons f-I curve when adaptation strength is increased [47]. In addition, SFA can be mediated by sodium current inactivation which cause the excited population of neurons to become less depolarized via the inactivation of Na+ inward flow [69]. That said, our SFA model appropriately captures their canonical role as slow, hyperpolarizing feedback mechanisms which activate in an up-state. H-current based adaptation is mediated by a different channel dynamics, causing dichotomous effects. The responsible hyperpolarization-activated cyclic nucleotide gated (HCN) channels cause K+ outward flow, all the while not preventing Na+ inward flow [70], and have a reversal potential between -50 and -20 mV [15]. Unlike potassium channels who have a clear selection of K+ (see [71]), HCN channels have a permeability of 3:1 to 5:1 for K+:Na+ from [15, 72, 73]. HCN channels are active as long as the neurons conductance is close to its resting membrane potential, causing positive inward currents that effectively excite the neuron. This, we simulate here, whereas the stagnating inhibitive effect for higher values of the membrane potential is neglegted similar to the cortical population model of [21]. Clearly, dynamical equivalence between SFA and h-current based adaptation is only expected to strictly hold for the approximate description of the adaptation dynamics given by Equations (1) and (4).
@@
Further studies could extend our investigations to a biophysically more realistic model, for example, as in [47], which distinguishes subthreshold activated versus spike-triggered SFA. [44] show for the model of [47] that subthreshold activated SFA changes the neuronal input-output relationship, and can increase the variability of interspike intervals (ISI) independent of current inputs, whereas spike-triggered SFA changes the ISIs variability based on the input to the neurons. It would be interesting to see, how far the diverse types of activations in SFA would break the shown symmetry to h-current based adaptation. For both adaptation mechanisms, the work of [22, 76] for non-spatial conductance based-models would allow to simulate other potentially symmetry-breaking channel dynamics, e.g., the stagnating inhibitive effect of h-currents. The Wilson-Cowan model investigated in this study may serve as a reference when exploring the effects of a biphysically realistic channel dynamics.

### v3

@@ IV Characterization of regular spatio-temporal patterns in Hopf- and Turing-unstable states
Figure 5: Properties of the spatio-temporal patterns emerging in Hopf- and Turing-unstable states for increasing h-current strength. A $k_{max}$ values (top row), dominant spatial frequency $f_{x}$ (middle row), and dominant temporal frequency $f_{t}$ (bottom row) for each location in state space. The figure shows slices of state space spanned by the external input $(I_{e},I_{e})\in[-2.5,2]\times[-3.5,1]$ for $b\in\{0,-0.5,-1.0\}$. Purple colors denote the corresponding feature values. Darker colors indicate higher values, see color bar. Yellow denotes states of no spatio-temporal activity, light green denotes states with irregular patterns with a standard deviation $r$ of the Kuramoto-order parameter of $r>10^{-2}$. Thin lines denote the boundaries of the dynamical regimes shown in Fig. 2A. Green boxes correspond to the panels shown in D. B Mean (dots) and variance (bars) of $k_{max},\ k_{0},\ f_{t},$ and $f_{x}$ for $b\in\{-0.25,-0.5,-0.75,-1\}$ for the states located in the Hopf-unstable, Turing-unstable up-, and Turing-unstable down-state regimes, for regular spatio-temporal activity for which the standard deviation of the Kuramoto-order parameter over time and space was $r\leq 10^{-2}$ (see Appendix B.3). C Histograms and kernel density estimates (KDE; thin lines) of the numerically calculated spatial frequency ($f_{x}$; top row), the spatial frequency corresponding to $k_{0}$ ($\frac{L\cdot k_{0}}{2\pi}$; middle row), and the spatial frequency corresponding to $k_{max}$ ($\frac{L\cdot k_{max}}{2\pi}$; bottom row) in the regimes of Hopf instability (left column), Turing instability in down-states (middle column), and Turing instability in up-states (right column) for b=0.5b=-0.5. Dashed vertical lines with numbers denote the location of the peak of the KDE. D Close-ups of state space (see green boxes in A) with values for $k_{max}$ (lower right), $f_{x}$ (upper right), and $f_{t}$ (upper left) in the Turing-unstable up- (left panels) and down-states (right panels) in bi. Yellow denotes states of no spatio-temporal activity or with irregular patterns with a standard deviation $r$ of the Kuramoto-order parameter of $r>10^{-2}$. The other colors denote feature values (see color bars). Wavenumbers were acquired as described in Section II.2; spatial and temporal frequencies as described in the Appendices B.1 and B.2. Numerical simulations were conducted with initialization close to the corresponding unstable fixed points.
@@
Figure 5 shows the results of the numerical investigation for the system with h-currents with increasing adaptation strength for the regimes of Hopf or Turing instability, where regular patterns occur. We did not consider the regimes of three fixed points (corresponding to dark and light beige in Fig. 2A), because numerical simulations close to the fixed points undergoing a Turing bifurcation lead to irregular propagation of activity fronts similar to the pattern exemplified in Fig. 2B, panel 7, which shows irregular activity emerging in a Turing-unstable up-state in bi.
@@
Figure 5A shows the wavenumber $k_{max}$ (top row), and the spatial, $f_{x}$ (center row) and temporal frequencies, $f_{t}$ (bottom row) of the emerging patterns for different locations in state space and for increasing adaptation strength. $k_{max}$ values vary continuously across state space. Values are larger than $0.44$ with highest values occurring in Turing-unstable states (see color bar). The spatial frequency, $f_{x}$, of the wave patterns change smoothly across state space for $b=0$ with the highest values occurring toward the center of the Hopf regime. Values decrease toward the borders. For b=0.5b=-0.5 and b=1b=-1, the distribution of $f_{x}$ becomes less smooth with non-equidistant, irregular patterns ($r>10^{-2}$) emerging toward the borders of the Hopf regime, where the regimes of Turing instability, one stable fixed point, Hopf instability, and three fixed points with only one stable state meet (green areas inside instability regimes, Fig. 5A). The numerical simulations also confirm that the activity traces for $b=0$ are stationary spatial waves, as $f_{t}$ is zero. For adaptation strong enough, waves start moving resulting in a finite value of $f_{t}$ which increases with the value of $|b|$ (see also Fig. 5B) confirming what is expected from the mathematical analysis. For increasing $|b|$, the averages of the wavenumbers $k_{0}$, and $k_{max}$ increase in value, albeit the variance over state space is high. $f_{x}$ increases with $|b|$ consistent with the increase of the wavenumbers $k_{0}$ and $k_{max}$. Additionally, the average temporal frequency increases steeply, which is expected from the increasing imaginary parts at the zero crossings of the real parts of the eigenvalues as discussed in Section III.
@@
Figure 6: Propagation speed of spatio-temporal patterns emerging in Hopf- and Turing-unstable states for increasing adaptation strength. A Slices of state space spanned by the external input currents showing speed values for the corresponding location in state space. Adaptation strength increases from left to right. Speed values are indicated by purple (see color bar), yellow denotes states of no spatio-temporal activity, green denotes states of irregular patterns with $r>10^{-2}$. Thin lines denote the borders of the dynamical regimes shown in Fig. 2A. B Histograms of the distribution of speed values across the states shown in A. Dashed vertical lines and numbers denote the location of the peaks of the histograms. C Average (dots) and variance (bars) of speed values across state space for $b\in\{-0.25,-0.5,-0.75,-1.0\}$. D spatio-temporal patterns of $u_{e}(x,t)$ for different values of $b$ emerging in the locations in state space which are indicated by hexagons in A. The corresponding value of the external input currents are given in Table 4 in Appendix B.5. The corresponding speed values are for 1: <<0.5; 2 and 3: 11.6; 4: 17.7; 5: 14.6; 6: 18.7. All other parameters are given in Table 1. E Close-ups of state spaces (see green boxes in Fig. 6A) with values for speed. Numerical simulations were initialized around the fixed points of Hopf or Turing instability. Speed values were computed as described in Appendix B.4.
In Fig. 5C we compare the spatial frequencies of the simulated activity patterns (top row) with the expected spatial frequencies corresponding to the two wavenumbers $k_{0}$, i.e., $\frac{L\cdot k_{0}}{2\pi}$ (middle row), and $k_{max}$, i.e., $\frac{L\cdot k_{max}}{2\pi}$ (bottom row). Since $k_{max}$ identifies the most rapidly growing unstable mode, expectations are to have highest similarity of spatial frequencies and frequencies corresponding to $k_{max}$. The distribution of spatial frequencies across Hopf states roughly agrees with this hypothesis (The Wasserstein distance between the distributions of $f_{x}$ and $\frac{L\cdot k_{max}}{2\pi}$ is $3.2$ and $12.9$ between the distributions of $f_{x}$ and $\frac{L\cdot k_{0}}{2\pi}$.), the spatial frequencies observed in Turing-unstable down- and up-states, however, agree better with $k_{0}$ (The Wasserstein distance between the distributions of $f_{x}$ and $\frac{L\cdot k_{max}}{2\pi}$ are larger than $14$ and smaller than $10$ between the distributions of $f_{x}$ and $\frac{L\cdot k_{0}}{2\pi}$.).
Moreover, spatial frequencies in Hopf-unstable states often lie in between $k_{0}$ and $k_{max}$, whereas for Turing-unstable states values are slightly below the spatial frequencies corresponding to $k_{0}$ which likewise held for simulated example activity traces in Hopf- and Turing-unstable states. These findings confirm that there is a range of spatial modes whose properties the activity patterns can inherit (called stability balloon in [63] [^6] In [63], a stability balloon describes a region in an $(R,\vec{q})$ plane, where $R$ is a control parameter to take the system further away from the equilibrium and $\vec{q}$ the wave vector effectively representing the wavenumber in a spatially one-dimensional ring.). Since we summarize fixed points stable at $k=0$ undergoing a Turing bifurcation at $k_{0}>0$ under Turing-unstable states, we know that $k_{0}$ is the smallest wavenumber at which the homogeneous equilibrium becomes unstable, i.e., determining the broadest wave. Nonetheless, we see that waves tend to converge to spatio-temporal patterns with a spatial frequency slightly below the expected frequency corresponding to the smallest wavenumber $k_{0}$, all the while the exact frequency varies depending on the initialization (see Appendix B.7, Fig. 11).
Since we observe bistability in Hopf-unstable states between spatially homogeneous oscillations and temporally stationary spatial oscillations (see Section III, we performed the same analysis as in Figures 5A and D in Appendix B.6 but initialized close to zero. Results are similar and shown in Appendix B.6 Fig. 10. This figure also shows the $k_{0}$ values over slices of state space corresponding to Fig. 5A.
Figure 5D shows close-ups of Turing unstable up- and down-states in bi for b=0.5b=-0.5. $k_{max}$ values do not deviate strongly from the $k_{max}$ values of the rest of the considered state spaces but only few numerical simulations converged to regular spatio-temporal patterns (non-yellow locations in the close-ups of state space). The resulting patterns strongly depend on the initialization, and even initializations close to the Turing-unstable fixed point do not necessarily converge to a spatio-temporal pattern but to the coexisting stable fixed point, as exemplified in Fig. 2B, panel 5(*), or irregular patterns [^7] States marked in yellow in the regions of Turing-unstable up- and down-states in bi shown in Fig. 5D and Fig. 6E may also support regular spatio-temporal patterns if another initialization would have been chosen., as exemplified in Fig. 2B, panel 7. If regular spatio-temporal patterns emerge, traveling waves show low values of spatial frequency. Temporal frequencies are also lower compared to the rest of state space.
@@
Figure 6B shows the distribution of speed values across Hopf-unstable and Turing-unstable states. The stronger the adaptation, the higher the speed values are. For $b=0$, as expected, activity patterns do not travel across space (recall Fig. 2, panel 0) or their speed is <0.5<0.5 (see Fig. 6D, panel 1). For b=0.5b=-0.5, there are two different speed values appearing the most often, with the dominant speed being $11.2$. For b=1b=-1, the variance of the distribution is larger and the most common speed value is once again higher at $17.6$. Figure 6C confirms that the average speed across state space increases with adaptation strength. Figure 6D shows traveling waves for the same relative locations in state space (i.e., considering the shift to the left discussed in Section III), increasing in adaptation strength from left to right column. “Broad" patterns become “narrower", increasing both temporal and spatial frequencies and their slope in the space-time diagrams, henceforth, the speed.
Figure 6E shows the same close-ups as Fig. 5D of Turing-unstable up- and down-states in bi showing the speed values for the numerical simulations converging to regular traveling waves. Speed values tend to be higher closer to the bifurcation line to the regime of single Turing-unstable up- and down-states.
@@ V Discussion
A semi-analytical stability analysis across parameter space was performed for increasing adaptation strength resulting in the identification of over ten dynamical regimes including two types of bistability. Spatio-temporal patterns, most often periodic adaptation-driven traveling waves, emerge in states exhibiting Hopf and Turing instabilities if adaptation is sufficiently strong. Activity patterns remain stationary in the absence of adaptation which is explained by the zero imaginary part of all eigenvalues of the linearization matrix when a fixed point undergoes a Turing bifurcation. This aligns with the results presented in [43], who prove the existence of stationary patterns for the reduced system if adaptation is sufficiently weak while traveling waves occur for stronger adaptation.
Changing the adaptation strength not only alters the model’s dynamical regimes but also the features of the emerging temporal and spatial patterns including the speed of traveling wave activity. The numerical simulations aligned with the mathematical analysis performed before. With stronger adaptation, wavenumbers increase on average, causing more waves on the ring, i.e., higher dominant spatial frequencies. Analogously, with growing imaginary parts of the eigenvalues at the zero crossing of their respective real parts for stronger adaptation, temporal frequencies also increase. These changes correlate with increasing traveling speeds of activity patterns when the adaptation becomes stronger.
@@
In our population-based model, both adaptation mechanisms were implemented in a simplified way. Hyperpolarizing feedback current of somatic SFA can be mediated by potassium channels, which activate when the excitatory neurons are in a depolarized high-activity state, effectively causing K+ outward flow [11, 65]. Different potassium channels have different properties leading to differential effects in the actual adaptation mechanism: Voltage- ($I_{m}$, see [66]) and sodium-activated K+-channels ($I_{KNa}$, see [67]) contribute mostly to subtreshold, while Ca2+-activated K+-channels ($I_{KCa}$, [68]) contribute to spike-based adaptation. Different channel types also have different effects on the response properties of a neuron, for example, on a neurons f-I curve when adaptation strength is increased [47]. In addition, SFA can be mediated by sodium current inactivation which cause the excited population of neurons to become less depolarized via the inactivation of Na+ inward flow [69]. That said, our SFA model appropriately captures their canonical role as slow, hyperpolarizing feedback mechanisms which activate in an up-state. H-current based adaptation is mediated by a different channel dynamics, causing dichotomous effects. The responsible hyperpolarization-activated cyclic nucleotide gated (HCN) channels cause K+ outward flow, all the while not preventing Na+ inward flow [70], and have a reversal potential between -50 and -20 mV [15]. Unlike potassium channels who have a clear selection of K+ (see [71]), HCN channels have a permeability of 3:1 to 5:1 for K+:Na+ [15, 72, 73]. HCN channels are active as long as the neurons conductance is close to its resting membrane potential, causing positive inward currents that effectively excite the neuron. This, we simulate here, whereas the stagnating inhibitive effect for higher values of the membrane potential is neglegted similar to the cortical population model of [21]. Clearly, dynamical equivalence between SFA and h-current based adaptation is only expected to strictly hold for the approximate description of the adaptation dynamics given by Eqs. (1) and (3).
@@
Further studies could extend our investigations to a biophysically more realistic model, for example, as in [47], which distinguishes subthreshold activated versus spike-triggered SFA. [44] show for the model of [47] that subthreshold activated SFA changes the neuronal input-output relationship, and can increase the variability of interspike intervals (ISI) independent of current inputs, whereas spike-triggered SFA changes the ISIs variability based on the input to the neurons. It would be interesting to see, how far the diverse types of activations in SFA would break the shown symmetry to h-current based adaptation. For both adaptation mechanisms, the work of [22, 76] for non-spatial conductance based-models would allow to simulate other potentially symmetry-breaking channel dynamics, e.g., the stagnating inhibitive effect of h-currents. The Wilson-Cowan model investigated in this study may serve as a reference when exploring the effects of a biophysically realistic channel dynamics.